What is isoelectric point? The amino acid glycine has two ionizable groups: 1) a carboxyl acid group with a pKa of 2.72 and 2) an amino group with a pKa of 9.60. Calculate the pI of glycine.
Answer #1

isoelectric point is nothing but a pH at which amino acid doent migrate which is called PI

Pi of the glycine = PKa1 + PKa2 / 2

PI = 2.72 + 9.60 / 2

PI = 12.32 / 2

PI = 6.16